OpenStudy (dls):

I need a small physics help since the physics section isn't loading for me..please help! A student says that he had applied force F=−k√x on a particle and the particcle moved in SHM..he refuses to tell whether k is a constnt or not..assume that he has worked only wth +ve x a xis and no othe r force acts on the particle..then? a) x increases as k increases b) x increases as k decreases c)x increases ,k=constant d)indeterminate

OpenStudy (dls):

attempt: \[\LARGE y=asinwt\] \[10=20\sin \frac{2\pi}{T}t\]

OpenStudy (anonymous):

144 [s]

OpenStudy (dls):

2s :P

OpenStudy (anonymous):

The mean position is 0cm.

OpenStudy (anonymous):

It's at equilibrium at 0cm since it is in the middle of the cycle

OpenStudy (anonymous):

Your equation is correct

OpenStudy (anonymous):

If you sub in T as 12

OpenStudy (anonymous):

then you find t

OpenStudy (dls):

will it be done like this? \[\LARGE \frac{1}{2}=tsin \frac{1}{2} \] t=1s so total=1*2=2s

OpenStudy (dls):

but which T as 12s? there are 2T's right,how are they different

OpenStudy (anonymous):

Big T

OpenStudy (anonymous):

Big T = 12

OpenStudy (dls):

what is the small t?

OpenStudy (anonymous):

t =1 if we plug it into Wolframalpha

OpenStudy (dls):

what does the small t represent?

OpenStudy (anonymous):

Lol Oops

OpenStudy (anonymous):

t is inside sine

OpenStudy (dls):

so what does it mean?

OpenStudy (anonymous):

small t represent the time and y represents where the particle is at time t

OpenStudy (dls):

what does big T represent :S

OpenStudy (anonymous):

T represents how long it take for the particle to get back to it original position

OpenStudy (anonymous):

So if we start at 0 cm, it will take 12 seconds to get back to 0 cm.

OpenStudy (dls):

THANKS! :D

OpenStudy (anonymous):

Got it? Here is the answer if you want to check it. http://www.wolframalpha.com/input/?i=t%3Dasin%281%2F2%29*6%2Fpi

OpenStudy (dls):

answer is 2s

OpenStudy (anonymous):

Yup. It's 2 seconds

OpenStudy (dls):

yes! twice of that.. can we try another question?

OpenStudy (anonymous):

Sure

OpenStudy (dls):

A particle under the action of SHM has a period of 3second and under the influence of another,it has a time period of 4 seconds..what will be its speed under the combined effect of the two?

OpenStudy (anonymous):

I'm not sure what the question means by "under the combined effects of the two".

OpenStudy (dls):

resultant of the two maybe

OpenStudy (anonymous):

The particle slows down since the particle's cycle slows down from 3s to 4s

OpenStudy (dls):

i have a formula here \[\LARGE T=\frac{T_{1} T_{2}}{\sqrt{T_{1}^{2}-T_{2}^{2}}} \]

OpenStudy (dls):

okay leave this one if its unclear !

OpenStudy (anonymous):

i don't recognize this equation sorry

OpenStudy (anonymous):

Perhaps you can open this question up on a new thread

OpenStudy (dls):

I dont want to spoil maths section by physics :S

OpenStudy (anonymous):

Physic is applied math

OpenStudy (anonymous):

lol

OpenStudy (dls):

:|

OpenStudy (dls):

2 pendulums have time period T and 5T/4 they start SHM at the same time from mean position..the phse difference between them after the bigger particle has completed 1 oscilation=?

OpenStudy (dls):

its pendulum and not particle*

OpenStudy (dls):

do u know this one?

OpenStudy (anonymous):

My knowledge in high school physics has limits. :|

OpenStudy (dls):

okay :o A student says that he had applied force \[\LARGE F=-k \sqrt{x}\] on a particle and the particcle moved in SHM..he refuses to tell whether k is a constnt or not..assume that he has worked only wth +ve x a xis and no othe r force acts on the particle..then?

OpenStudy (dls):

a) x increases as k increases b) x increases as k decreases c)x increases ,k=constant d)indeterminate

OpenStudy (anonymous):

\[y=a \sin 2Pift\] a is amplitude =20cm, f=1/T=1/12 hz, y is the displacement, = 20cm (10cm on each side) subst into the equation \[20=20\sin \pi \times 2\times t \div12\] \[\pi \times t \div 6= \sin^{-1} 1\] \[\pi t = 1.57 \times 6 = 9.42\] \[t=9.42\div3.14=3\] t=3seconds

OpenStudy (anonymous):

For the last question, imagine this. You are holding a spring on one hand with your left hand and you are pulling the other end with your right. As you pull further right, you feel a strong force on your right hand to the left. Now let's apply math. Let say x = 0 stands for the spring at equilibrium. As you pull the to the right, x increases positively and remember k is named "the spring constant".

OpenStudy (dls):

okay

OpenStudy (anonymous):

Notice the -k in the equation. There is a negative because the string wants to restore its position whenever it is stretched or compressed.

OpenStudy (dls):

yes

OpenStudy (anonymous):

my answer would be c)x increases ,k=constant

OpenStudy (dls):

its a :P

OpenStudy (anonymous):

k changes?

OpenStudy (dls):

yes..how

OpenStudy (anonymous):

Is this answer from your textbook?

OpenStudy (dls):

yes

OpenStudy (anonymous):

From wikipedia, k is the rate, spring constant or force constant of the spring, a constant that depends on the spring's material and construction

OpenStudy (anonymous):

Unless the spring changes its material or its construction, I don't think k will change.

OpenStudy (dls):

ITS not given that k-constant read the ques

OpenStudy (anonymous):

Hooke's law is \(F=-kx\) not \(-k\sqrt{x}\)

OpenStudy (dls):

but he assumed thats given in question

OpenStudy (dls):

we dont have to generalize anything i guess..like we dont even know

OpenStudy (dls):

if k is a constant or not

OpenStudy (unklerhaukus):

check the units

OpenStudy (anonymous):

If it's undergoing simple harmonic motion with a single force, that means \(F\propto x\)... i.e. \(-k\propto\sqrt{x}\), so that the force can be written \(F=-k\sqrt{x}=-Kx\) where \(k=K\sqrt{x}\) and therefore non-constant. This is consistent with \(x>0\), so we stay in the reals. It should be clear that since \(k=K\sqrt{x}\), and \(K\) is positive (as the spring constant in \(F=-Kx\)), that an increase in \(x\) translates to an increase in \(k\).

OpenStudy (dls):

eh?

OpenStudy (dls):

simple language?

OpenStudy (anonymous):

If it's undergoing simple harmonic motion, according to Hooke's law the force must be directly proportional to the displacement i.e. \(F=-Kx\) where \(K\) is our spring constant (always positive). The only force here is \(F=-k\sqrt{x}\), which doesn't immediately resemble our above equation. If we set \(-k\sqrt{x}=-Kx\), you can solve for \(k\) to yield \(k=K\sqrt{x}\). Since \(K\) is positive, and \(x\) is positive, we can see that an increase of \(x\) will cause in an increase in \(\sqrt{x}\) and thus in \(k\).

OpenStudy (dls):

thanks!